verify that x=-2 is solution of equation 3(2x+3) - x*x = 2x + 7
Answer by RAY100(1637)

Normal method to verify is to substitute answer into eqn
.
3(2x+3) -x*x = 2x +7
.
3 [ 2(-2) +3] -(-2)*(-2) = 2(-2) +7
.
3[ -1] -4 = -4+7
.
-3-4=3
.
-7 = 3,,,,not true, not verified
.
lets just solve eqn
.
3(2x+3) -x*x =2x +7
.
6x +9 -x^2 = 2x+7
.
0 = x^2 -4x -2
.
use quadratic eqn to solve,,,,,a=1,,,,b=-4,,,,,c=-2
.
x = [ -(-4) +/- sqrt{ (-4)^2 -4 (1) (-2)} ] / 2(1)
.
x=[ 4 +/- sq rt { 16 +8} ] /2
.
x= [ 4 +/- sqrt 24]/2
.
x= [ 4 +/- 4.9] /2
.
x= +4.4, -.45
.
to verify these
.
x^2 -4x -2 =0
.
(4.4)^2 -4(4.4) -2 =0,,,,,ok
.
(.45)^2 -4(.45) -2 =0,,,,,ok